OpenStudy (anonymous):

Order fractions least to greatest 5/8 2/3 3/5

OpenStudy (unklerhaukus):

convert your fractions to a common base,

OpenStudy (anonymous):

try converting into decimals 5/8=0.625 2/3=0.66666666666666666666666666 3/5=0.6 so the order would be 3/5, 5/8, 2/3 note: You should really learn this stuff. I knew it in 4th grade, and I am in 7th grade now.

OpenStudy (mertsj):

\[\frac{5}{8}=\frac{75}{120}\] \[\frac{2}{3}=\frac{80}{120}\] \[\frac{3}{5}=\frac{72}{120}\]

OpenStudy (anonymous):

I know it but I'm lazy.

OpenStudy (unklerhaukus):

Mertsj has found a common base 120, this show all the fractions on the same scale you can now simply compare the numerators